the length of one leg is 6ft. The length of the hypotenuse is 3ft longer than the other leg. Find the length of the hypotenuse and the other leg.
Answer by nerdybill(7384)

the length of one leg is 6ft. The length of the hypotenuse is 3ft longer than the other leg. Find the length of the hypotenuse and the other leg.
.
Let x = length of "other leg"
then
x+3 = length of hypotenuse
.
applying Pythagorean theorem:
x^2 + 6^2 = (x+3)^2
x^2 + 6^2 = (x+3)(x+3)
x^2 + 36 = x^2+6x+9
36 = 6x+9
27 = 6x
27/6 = x
4.5 feet = x (other leg)
.
hypotenuse:
x+3 = 4.5+3 = 7.5 feet